I need help please solve for V!!
8VW=33-V

Answer by lynnlo(4176)

8VW=33-V
WRITE THE LINEAR POLYNOMIAL ON THE LEFT SIDE IN STANDARD FORM
COLLECT IN TERMS OF V
V(8W+1)=33
DIVIDE BOTH SIDES BY 8W+1
V=33/8W+1